The firm is Costco Wholesale Corporation. Beta is 0.67. Please provide the following

a. A SML calculation for your firm’s stock, using this beta and 5 year averages of historical market (NYSE or S&P 500) returns and historical Treasury bill returns.

b. A comparison of the SML required return on your firm’s stock and the discounted stock valuation required return on the stock. {P0= D1/ (rs-g) implies rs = D1/P0 + g}.

1
Expert's answer
2021-01-26T03:17:47-0500

a. Expected rate of return:

"E(R_{cost}) = R_F + \u03b2[E(R_M) \u2013 R_F]"

"R_F" is rate of return on historical Treasury bill returns

"\u03b2" is systematic risk

"E(R_M)" is expected rate of return on market portfolio (5 year averages of historical market, S&P 500, 2017 - 2021)

"E(R_M)=\frac{19.42-6.24+28.88+16.26+2.58}{5}=12.18" %

"E(R_{cost})=1.98+0.67(12.18-1.98)=8.81" %


b. Rate of Return = (Dividend Payment / Stock Price) + Dividend Growth Rate

Dividend Payment"=2.75" %

Stock Price"=9.02" %

Dividend Growth Rate"=12.59" %

Required Rate of Return"=2.75\/9.02+12.59=12.89" %


So, we have:

SML Rate of Return"<" Discounted Stock Valuation Return